An airplane is traveling NE, 320 miles an hour due north, and 40 miles an hour due east.
The y axis is north, the x axis is east. How far will the plane have traveled after 15 minutes?

Answer by stanbon(75887)

An airplane is traveling NE, 320 miles an hour due north, and 40 miles an hour due east.
The y axis is north, the x axis is east. How far will the plane have traveled after 15 minutes (1/4 hour) ?
---------------------
Draw the picture:
You have a right triangle with height = (1/4)(320) = 80 miles,
and base = (1/4)(40) = 10 miles
------------
Ans: Distance = sqrt[80^2 + 10^2] = sqrt(6500) = 80.62 miles
--------------
Cheers,
Stan H.
--------------